OpenStudy (anonymous):

What is the area of a rhombus with a side of 9 and an altitude of 6?

OpenStudy (jiteshmeghwal9):

since the area of rhombus = \(\Large{\frac{1}{2}\times base \times altitude}\) therefore by putting the values we have, \(\Large{\frac{1}{2}\times9\times6}\) \(\Large{9\times3=27}\)

OpenStudy (anonymous):

thank you

OpenStudy (jiteshmeghwal9):

yw :)